OpenStudy (somy):

number of moles after equilibrium has reached please explain this to me, i have no clue what to think of

OpenStudy (somy):

@emcrazy14 @Kainui

OpenStudy (somy):

i don't know what to do, im not even given volume to find concentration or at least there is no mention of at room temperature and pressure or at standard temperature and pressure T_T

OpenStudy (anonymous):

You have to do this by the equilibrium (Kc) method.

OpenStudy (anonymous):

@somy: My bad. That's actually Kc= x^2/(1-x)^2

OpenStudy (anonymous):

\[4=\frac{ x ^{2} }{ (1-x)^{2} }\]

OpenStudy (kainui):

@Somy they didn't give you volume, but it's actually not important and I'll show you why! Since equilibrium constants are concentrations, and concentration is recorded as mol/L then let's just plug in concentrations as capital letters with brackets and the number of mols as the lower case letter with V as the volume. \[K_c=\frac{[C][D]}{[A][B]}=\frac{ \frac{c}{V} \frac{d}{V}}{ \frac{a}{V} \frac{b}{V}}=\frac{cd}{ab}\] So you can use mols just as fine as concentrations, it all works out. =)

OpenStudy (somy):

ohhhhh okaaay let me try now @Kainui

OpenStudy (anonymous):

Taking square root on both sides; \[\sqrt{4}= \sqrt{\frac{ x ^{2} }{ (1-x)^{2} }}\] \[2= \frac{ x }{ (1-x) }\] \[2-2x=x\] \[3x=2\] \[x=\frac{ 2 }{ 3 }\]

OpenStudy (kainui):

@emcrazy14 showed the correct technique of how to do it, I just wanted to give you a little extra info

OpenStudy (somy):

@emcrazy14 i'll try it myself then look at ur working ok?

OpenStudy (anonymous):

@somy: I'm sorry. Sure. I did that for @KL-RC . :)

OpenStudy (somy):

i do know that way of working, i was just confused about volume not being there :)

OpenStudy (somy):

ah no need to apologize lol i said that because me doing on my own might take time, and so i might have questions to ask :D so don't leave lmao @emcrazy14

OpenStudy (anonymous):

Okay. :D

OpenStudy (somy):

yuppp i got it right :D thank u guys!!! :* @Kainui @emcrazy14

OpenStudy (anonymous):

Well done. *_*

OpenStudy (somy):

how's physics? @emcrazy14

OpenStudy (anonymous):

Really bad. I get so many wrong. >_<

OpenStudy (somy):

im freaking out right now, tomorrow we are going to start with physics @emcrazy14

OpenStudy (somy):

which year r u doing? @emcrazy14

OpenStudy (anonymous):

Before 2005.

OpenStudy (somy):

O_O isn't that too far? @emcrazy14

OpenStudy (anonymous):

They rarely repeat the questions. So practising whatever year won't make a difference I guess. They will always come up with something new. >_<

OpenStudy (somy):

well i guess so also, but i thought latest years would be better T_T